# Delzant type theorem for torus-equivariantly embedded toric hypersurfaces

### v1

@@ Abstract.
In the previous work, we study the closure of a complex subtorus in a toric manifold given by the data of an affine subspace. We call it torus-equivariantly embedded toric manifold when the closure of a complex subtorus is a smooth complex submanifold. In this paper, we clearify the condition for nonsingularity of the closure of a complex subtorus in terms of polytopes. The main result is a generalization of Delzant theorem to the case of torus-equivariantly embedded toric hypersurfaces.
@@ 1. Introduction
Symplectic toric manifolds are $2n$-dimensional symplectic manifolds equipped with the effective Hamiltonian action of an $n$-dimensional torus. The image of the moment map for the Hamiltonian torus action is the convex hull of the image of the fixed points of the Hamiltonian action [Ati82, GS82]. Due to the work of Delzant [Del88], there is a one-to-one correspondence between symplectic toric manifolds and certain convex polytopes appeared as the moment polytopes of symplectic toric manifolds. The moment polytopes of symplectic toric manifolds are called Delzant polytopes. Moreover, symplectic toric manifolds canonnically admit a Kähler metric, which is called a Guillemin metric [Gui94a, Gui94b, CDG03]. Using the data of a Delzant polytope, the complements of the toric diviors $D$ in the corresponding symplectic toric manifold $X$ can be identified with the complex torus $(\mathbb{C}^{\ast})^{n}$.
@@
In [Yam24b], we construct the closure $\overline{C(V)}$ of the $k$-dimensional complex subtorus $C(V)$ in the toric divior complement $X\setminus D\cong(\mathbb{C}^{\ast})^{n}$ from the data of a $k$-dimensional affine subspace $V\subset\mathfrak{t}^{n}\cong\mathbb{R}^{n}$. The closure $\overline{C(V)}$ can be expressed by the zero locus of polynomials $f^{\lambda}_{k+1},\ldots,f^{\lambda}_{n}$ for each vertex $\lambda$ of the Delzant polytope, which are determined by the data of the Delzant polytope and the affine subspace. Note that the closure $\overline{C(V)}$ might have singularity at the intersection with the toric diviors $D$. We show in [Yam24b, Theorem 4.20] that if the closure $\overline{C(V)}$ is a smooth $k$-dimensional complex submanifold in the symplectic toric manifold $X$, then the moment polytope for the Hamiltonian $k$-dimensional torus action on $\overline{C(V)}$ coincides with the moment polytope for the Hamiltonian $k$-dimensional subtorus action on $X$. Moreover, the submanifold $\overline{C(V)}$ is a symplectic toric manifold with respect to the $k$-dimensional torus action on $\overline{C(V)}$. In particular, if $\overline{C(V)}$ is a complex submanifold, then the moment polytope for the Hamiltonian $k$-dimensional subtorus action on $X$ is a Delzant polytope. We call the complex submanifold $\overline{C(V)}$ the torus-equivariantly embedded toric manifold.
@@
Even though the moment polytope for the Hamiltonian $k$-dimensional subtorus action on $X$ is a Delzant polytope as a convex polytope for some affine subspace $V$, $\overline{C(V)}$ might not be a smooth complex submanifold. For example, let $X=\mathbb{C}P^{2}$. For any $1$-dimensional affine subspace $V$ in $\mathfrak{t}^{2}\cong\mathbb{R}^{2}$, the moment polytope for the Hamiltonian one-dimensional subtorus action associated with $V$ becomes an interval in $(\mathfrak{t}^{1})^{\ast}\cong\mathbb{R}$, which can be seen as the Delzant polytope of $\mathbb{C}P^{1}$. However, there are many affine subspaces $V$ such that the closure $\overline{C(V)}$ has a singularity (see [Yam24b, Section 5.1]).
@@ 1.1. Main Results
In this paper, we show that a Delzant type correspondence for torus-equivariantly embedded toric manifolds of codimension one, i.e. $k=n-1$.
Let $V$ be an $(n-1)$-dimensional affine subspace in $\mathfrak{t}^{n}\cong\mathbb{R}^{n}$ with rational slope, i.e. $\overline{V}\cap\mathbb{Z}^{n}\cong\mathbb{Z}^{n-1}$, where $\overline{V}$ is the linear part of $V$. Then, we may take a primitive $\mathbb{Z}$-basis $p_{1},\ldots,p_{n-1}\in\mathbb{Z}^{n}$ of $\overline{V}\cap\mathbb{Z}^{n}$ and $a\in\mathbb{R}^{n}$ such that $V=\mathbb{R}p_{1}+\cdots+\mathbb{R}p_{n-1}+a$. We define the injective homomorphism $i_{V}:T^{n-1}\to T^{n}$ by
@@
Let $\Delta$ be a Delzant polytope and $V=\mathbb{R}p_{1}+\cdots+\mathbb{R}p_{n-1}+a$ an $(n-1)$-dimensional affine subspace in $\mathfrak{t}^{n}$ with rational slope. A vertex $\lambda$ of the polytope $\Delta$ is a good vertex with respect to the pullback $i_{V}^{\ast}:(\mathfrak{t}^{n})^{\ast}\to(\mathfrak{t}^{n-1})^{\ast}$ if the vertex $\lambda$ satisfies the two conditions:
@@
Let $V=\mathbb{R}p_{1}+\cdots+\mathbb{R}p_{n-1}+a$ be an $(n-1)$-dimensional affine subspace in $\mathfrak{t}^{n}$ with rational slope. A Delzant polytope $\Delta$ is a good polytope with respect to the map $i_{V}^{\ast}$ if any good vertex $\lambda$ with respect to the map $i_{V}^{\ast}$ satisfies the two conditions:
  • we can choose $n-1$ direction vectors $v^{\lambda}_{1},\ldots,v^{\lambda}_{j-1},v^{\lambda}_{j+1},\ldots,v^{\lambda}_{n}$ such that $i_{V}^{\ast}(v^{\lambda}_{1}),\ldots,i_{V}^{\ast}(v^{\lambda}_{j-1}),i_{V}^{\ast}(v^{\lambda}_{j+1}),\ldots,i_{V}^{\ast}(v^{\lambda}_{n})$ are linearly independent and $i_{V}^{\ast}(\{\sum_{i=1}^{n}a_{i}v^{\lambda}_{i}\mid a_{i}\geq 0\})=\{\sum_{i\neq j}b_{i}i_{V}^{\ast}(v^{\lambda}_{i})\mid b_{i}\geq 0\}$,
  • the direction vectors $i_{V}^{\ast}(v^{\lambda}_{1}),\ldots,i_{V}^{\ast}(v^{\lambda}_{j-1}),i_{V}^{\ast}(v^{\lambda}_{j+1}),\ldots,i_{V}^{\ast}(v^{\lambda}_{n})\in\mathbb{Z}^{n-1}$ from the vertex $i_{V}^{\ast}(\lambda)$ of $i_{V}^{\ast}(\Delta)$ form a $\mathbb{Z}$-basis of $\mathbb{Z}^{n-1}$.
Note that the first condition in Definition 1.2 are equivalent to the condition that the polytope $i_{V}^{\ast}(\Delta)$ is simple.
@@
Let $\Delta$ be a Delzant polytope and $V=\mathbb{R}p_{1}+\cdots+\mathbb{R}p_{n-1}+a$ an $(n-1)$-dimensional affine subspace in $\mathfrak{t}^{n}$ with rational slope. If the Delzant polytope $\Delta$ is good with respect to the map $i_{V}^{\ast}$, then the rank of the Jacobian matrix $Df^{\lambda}$ is equal to one at any point of the zero locus of $f^{\lambda}$ for any vertex $\lambda$ of the Delzant polytope $\Delta$. In particular, $\overline{C(V)}$ is a smooth complex hypersurface in the toric manifold $X$.
@@
If the Delzant polytope $\Delta$ of a toric manifold $X$ is good with respect to the map $i_{V}^{\ast}$, then the convex polytope $i_{V}^{\ast}(\Delta)$ is a Delzant polytope of the complex hypersurface $\overline{C(V)}$ in $X$.
@@
Let $V=\mathbb{R}p_{1}+\cdots+\mathbb{R}p_{n-1}+a$ be an $(n-1)$-dimensional affine subspace in $\mathfrak{t}^{n}$ with rational slope. If the rank of the Jacobian matrix $Df^{\lambda}$ is equal to one at any point of the zero locus of $f^{\lambda}$ for any vertex $\lambda$ of the Delzant polytope $\Delta$ of the toric manifold $X$, then the Delzant polytope $\Delta$ is good with respect to the map $i_{V}^{\ast}$.
@@ 1.3. Outline
This paper is organized as follows. In Section 2, we prepare for something about torus-equivariantly embedded toric hypersurfaces. In Section 3, we show Theorem 1.3. In Section 4, we show Theorem 1.5.
@@
The author is grateful to the advisor, Manabu Akaho, for a lot of encouragements and supports. The author would also like to thank Yuichi Kuno and Yasuhito Nakajima for helpful discussions. This work was supported by JST SPRING, Grant Number JPMJSP2156.
2. Reveiw of Torus-equivariantly Embedded Toric Manifolds
In this section, we review the settings in [Yam24b]. We also show some properties of the rank of the Jacobian matrix $Df^{\lambda}$ used in this paper.
@@ 2.1. Symplectic Toric Manifolds
[${{}^{t}[u^{\lambda}_{1}\;\cdots\;u^{\lambda}_{n}]}[v^{\lambda}_{1}\;\cdots\;v^{\lambda}_{n}]=E_{n}$
holds for any vertex $\lambda$ of $\Delta$ (see [Yam24b, Lemma 3.10] for example).
@@
The detailed construction in terms of the notation used here is given in [Yam24b, Section 2.1 and Section 3.2]. Moreover, the complement of toric diviors in a $2n$-dimensional symplectic toric manifold can be identified with a complex $n$-dimensional torus.
@@ 2.2. Torus-equivariantly Embedded Toric Hypersurfaces
Let $p_{1},\ldots,p_{n-1}\in\mathbb{Z}^{n}$ be primitive vectors which are linearly independent, and $a\in\mathbb{R}^{n}$. We consider the $(n-1)$-dimensional affine subspace $V=\mathbb{R}p_{1}+\cdots+\mathbb{R}p_{n-1}+a$ in $\mathfrak{t}^{n}\cong\mathbb{R}^{n}$. Write $\overline{V}$ as the linear part of $V$. Assume that $V$ has a rational slope, i.e. we may assume that $p_{1},\ldots,p_{n-1}\in\mathbb{Z}^{n}$ form a $\mathbb{Z}$-basis of $\overline{V}\cap\mathbb{Z}^{n}\cong\mathbb{Z}^{n-1}$. Let $q\in\mathbb{Z}^{n}$ be a primitive basis of the orthogonal subspace to $\overline{V}$ in $\mathfrak{t}^{n}\cong\mathbb{R}^{n}$.
Through the log-affine coordinate $(\mathbb{C}^{n})^{\ast}\cong\mathfrak{t}^{n}\times T^{n}$, we define the complex subtorus $C(V)\cong V\times T^{n-1}$ (see [Yam24b, Proposition 4.2] for detail). Recall that the complement of toric diviors in a toric manifold $X$ can be identified with a complex $n$-dimensional torus $(\mathbb{C}^{\ast})^{n}$. The closure $\overline{C(V)}$ of $C(V)$ in the toric manifold $X$ is expressed as follows:
@@
The closure $\overline{C(V)}$ of $C(V)$ in the toric manifold $X$ is defined as $\overline{C(V)}=\bigcup_{\lambda\in\Lambda}\varphi_{\lambda}^{-1}(\overline{C_{\lambda}(V)})$, where $\overline{C_{\lambda}(V)}\subset\varphi_{\lambda}(U_{\lambda})\cong\mathbb{C}^{n}$ is a zero locus of $f^{\lambda}$. Here, $f^{\lambda}$ are defined by
@@
The rank of the Jacobian matrix $Df^{\lambda}$ of $f^{\lambda}$ might be equal to zero at some points in the toric diviors. In particular, $\overline{C(V)}$ might have a singularity at some points in the toric diviors.
@@
Proof.
We show that the $T^{n-1}$-action on $\overline{C(V)}$ is effective, Hamiltonian. Define the injective homomorphism $i_{V}:T^{n-1}\to T^{n}$ by
$i_{V}(t_{1},\ldots,t_{n-1})=\left(\prod_{l=1}^{n-1}t_{l}^{\langle p_{l},e_{1}\rangle},\ldots,\prod_{l=1}^{n-1}t_{l}^{\langle p_{l},e_{n}\rangle}\right).$
Since the toric divior complement $X\setminus D$ can be identified with a complex torus $(\mathbb{C}^{\ast})^{n}$, we obtain $X\setminus D\cong\mathfrak{t}^{n}\times T^{n}$. Since $C(V)\cong(\mathbb{C}^{\ast})^{n-1}$ is a complex subtorus in $(\mathbb{C}^{\ast})^{n}\cong X\setminus D$, we obtain the identification $C(V)\cong V\times i_{V}(T^{n-1})\subset\mathfrak{t}^{n}\times T^{n}\cong X\setminus D$. Since we defined the $T^{n-1}$-action on $\overline{C(V)}$ through the map $i_{V}:T^{n-1}\to T^{n}$ and the $T^{n}$-action on $X$ (see [Yam24b]), we see that the $T^{n-1}$-action on $\overline{C(V)}$ is effective.
If $\overline{C(V)}$ is a complex submanifold in $X$, then there exists an embedding $i:\overline{C(V)}\to X$. Thus, a symplectic form on $\overline{C(V)}$ is $i^{\ast}\omega$, where $\omega$ is a symplectic form on $X$. Let $R_{t}:\overline{C(V)}\to\overline{C(V)}$ be a map defined by
$R_{t}(x)=x\cdot t,$
where $\cdot$ is the $T^{n-1}$-action on $\overline{C(V)}$. By the definition of the $T^{n-1}$-action on $\overline{C(V)}$, we obtain
$R_{t}^{\ast}(i^{\ast}\omega)=i^{\ast}R_{t}^{\ast}\omega=i^{\ast}\omega,$
for any $t\in T^{n-1}$. Thus, the $T^{n-1}$-action on $\overline{C(V)}$ preserves the symplectic form $i^{\ast}\omega$ on $\overline{C(V)}$. We show that the map $\overline{\mu}:=i_{V}^{\ast}\circ\mu\circ i:\overline{C(V)}\to(\mathfrak{t}^{n-1})^{\ast}$ is the moment map for the $T^{n-1}$-action on $\overline{C(V)}$, where $\mu:X\to(\mathfrak{t}^{n})^{\ast}$ is the moment map for the $T^{n}$-action on $X$. By the definition of the $T^{n-1}$-action on $\overline{C(V)}$, the map $\overline{\mu}$ is equivariant with respect to the $T^{n-1}$-action on $\overline{C(V)}$. For any point $p\in\overline{C(V)}$ and $X\in\mathfrak{t}^{n-1}$, we have
$\overline{\mu}_{X}(p)$ $=\langle i_{V}^{\ast}\mu(i(p)),X\rangle$
$=\langle\mu(i(p)),(i_{V})_{\ast}X\rangle$
$=\langle\mu(i(p)),X\rangle$
$=(\mu_{X}\circ i)(p).$
From this calculation, we obtain
$d\overline{\mu}_{X}$ $=d(\mu_{X}\circ i)$
$=di^{\ast}\mu_{X}$
$=i^{\ast}d\mu_{X}$
$=i^{\ast}\omega(X^{\sharp},-).$
Therefore, the map $\overline{\mu}$ is the moment map for the $T^{n-1}$-action on $\overline{C(V)}$ equipped with the symplectic form $i^{\ast}\omega$. ∎
Note that the above discussion is valid for any codimension of $\overline{C(V)}$.
@@
In this case, the rank of the Jacobian matrix $Df^{\lambda}$ is independent of the choice of $a\in\mathbb{R}^{n}$ of the affine subspace $V=\mathbb{R}p_{1}+\cdots+\mathbb{R}p_{n-1}+a$ [Yam24a, Proposition 2.12], i.e. we may assume that $a=0$.

### v2

@@ Abstract.
In the previous work, we study the moment polytope of the closure of the complex subtorus orbit in a symplectic toric manifold associated to an affine subspace when the closure is a smooth complex submanifold. In this paper, we clarify the condition for nonsingularity of the closure of the codimension one complex subtorus orbit in terms of polytopes. The main result is a generalization of Delzant theorem.
@@ 1. Introduction
Symplectic toric manifolds are $2n$-dimensional symplectic manifolds equipped with an effective Hamiltonian action of an $n$-dimensional torus. The image of the moment map for the Hamiltonian torus action is the convex hull of the image of the fixed points of the Hamiltonian action [Ati82, GS82]. Due to the work of Delzant [Del88], there is a one-to-one correspondence between symplectic toric manifolds and certain convex polytopes appeared as the moment polytopes of symplectic toric manifolds. The moment polytopes of symplectic toric manifolds are called Delzant polytopes. Moreover, symplectic toric manifolds canonically admit a Kähler metric, which is called the Guillemin metric [Gui94a, Gui94b, CDG03]. Using the data of a Delzant polytope, the complements of the toric divisors $D$ in the corresponding symplectic toric manifold $X$ can be identified with the complex torus $(\mathbb{C}^{\ast})^{n}$.
@@
In [Yam25], we construct the closure $\overline{C(V)}\subset X$ of the $(\mathbb{C}^{\ast})^{k}$-orbit $C(V)\subset X\setminus D\cong(\mathbb{C}^{\ast})^{n}$ from the data of a $k$-dimensional affine subspace $V\subset\mathfrak{t}^{n}\cong\mathbb{R}^{n}$. The construction of $\overline{C(V)}$ is motivated by Yamamoto’s construction [Yam18] of the Leung–Yau–Zaslow correspondence [LYZ00] in the semi-flat case of Strominger–Yau–Zaslow mirror symmetry [SYZ96]. The closure $\overline{C(V)}$ can be expressed by the zero locus of polynomials $f^{\lambda}_{k+1},\ldots,f^{\lambda}_{n}$ for each vertex $\lambda$ of the Delzant polytope, which are determined by the data of the Delzant polytope and the affine subspace. Note that the closure $\overline{C(V)}$ might have singularity at the intersection with the toric divisors $D$. We show in [Yam25, Theorem 4.20] that if the closure $\overline{C(V)}$ is a smooth $k$-dimensional complex submanifold in the symplectic toric manifold $X$, then the moment polytope for the Hamiltonian $k$-dimensional torus action on $\overline{C(V)}$ coincides with the moment polytope for the Hamiltonian $k$-dimensional subtorus action on $X$. Moreover, the submanifold $\overline{C(V)}$ is a symplectic toric manifold with respect to the $k$-dimensional torus action on $\overline{C(V)}$. In particular, if $\overline{C(V)}$ is a complex submanifold, then the moment polytope for the Hamiltonian $k$-dimensional subtorus action on $X$ is a Delzant polytope. We call the complex submanifold $\overline{C(V)}$ the torus-equivariantly embedded toric manifold.
@@
Even though the moment polytope for the Hamiltonian $k$-dimensional subtorus action on $X$ is a Delzant polytope as a convex polytope for some affine subspace $V$, $\overline{C(V)}$ might not be a smooth complex submanifold. For example, let $X=\mathbb{C}P^{2}$. For any $1$-dimensional affine subspace $V$ in $\mathfrak{t}^{2}\cong\mathbb{R}^{2}$, the moment polytope for the Hamiltonian one-dimensional subtorus action associated with $V$ becomes an interval in $(\mathfrak{t}^{1})^{\ast}\cong\mathbb{R}$, which can be seen as the Delzant polytope of $\mathbb{C}P^{1}$. However, there are many affine subspaces $V$ such that the closure $\overline{C(V)}$ has a singularity (see [Yam25, Section 5.1]).
Under the Delzant correspondence, we expect that objects in symplectic toric geometry shall be characterized combinatorially in terms of polytopes. The main result in this paper is to give a characterization of the nonexistence of singular points in $\overline{C(V)}$ in terms of (moment) polytopes. Moreover, through the Leung–Yau–Zaslow correspondence, we expect that this characterization must have a relation with an asymptotic condition of conormal bundles of $V$ in $\mathfrak{t}^{n}$ (in the sense of [Cha09, Han19] for example).
@@ 1.1. Main Results
In this paper, we show that a Delzant type correspondence for torus-equivariantly embedded toric manifolds of codimension one, i.e. k=n1k=n-1.
Let $V$ be an (n1)(n-1)-dimensional affine subspace in $\mathfrak{t}^{n}\cong\mathbb{R}^{n}$ with rational slope, i.e. $\vec{V}\cap\mathbb{Z}^{n}\cong\mathbb{Z}^{n-1}$, where $\vec{V}$ is the linear part of $V$. Then, we may take a primitive $\mathbb{Z}$-basis $p_{1},\ldots,p_{n-1}\in\mathbb{Z}^{n}$ of $\vec{V}\cap\mathbb{Z}^{n}$ and $a\in\mathbb{R}^{n}$ such that $V=\mathbb{R}p_{1}+\cdots+\mathbb{R}p_{n-1}+a$. We define the injective homomorphism $i_{V}:T^{n-1}\to T^{n}$ by
@@
Let $\Delta$ be a Delzant polytope in $(\mathfrak{t}^{n})^{\ast}$ and $V=\mathbb{R}p_{1}+\cdots+\mathbb{R}p_{n-1}+a$ an (n1)(n-1)-dimensional affine subspace in $\mathfrak{t}^{n}$ with rational slope. A vertex $\lambda$ of the polytope $\Delta$ is a good vertex with respect to the pullback $i_{V}^{\ast}:(\mathfrak{t}^{n})^{\ast}\to(\mathfrak{t}^{n-1})^{\ast}$ if the vertex $\lambda$ satisfies the two conditions:
@@
Let $\Delta$ be a Delzant polytope in $(\mathfrak{t}^{n})^{\ast}$. An affine subspace $V$ in $\mathfrak{t}^{n}$ with rational slope is admissible with respect to the Delzant polytope $\Delta$ if any good vertex $\lambda$ with respect to the map $i_{V}^{\ast}$ satisfies the two conditions:
  • we can choose n1n-1 direction vectors $v^{\lambda}_{1},\ldots,v^{\lambda}_{j-1},v^{\lambda}_{j+1},\ldots,v^{\lambda}_{n}$ such that $i_{V}^{\ast}(v^{\lambda}_{1}),\ldots,i_{V}^{\ast}(v^{\lambda}_{j-1}),i_{V}^{\ast}(v^{\lambda}_{j+1}),\ldots,i_{V}^{\ast}(v^{\lambda}_{n})$ are linearly independent and $i_{V}^{\ast}(\{\sum_{i=1}^{n}a_{i}v^{\lambda}_{i}\mid a_{i}\geq 0\})=\{\sum_{i\neq j}b_{i}i_{V}^{\ast}(v^{\lambda}_{i})\mid b_{i}\geq 0\}$,
  • the vectors $i_{V}^{\ast}(v^{\lambda}_{1}),\ldots,i_{V}^{\ast}(v^{\lambda}_{j-1}),i_{V}^{\ast}(v^{\lambda}_{j+1}),\ldots,i_{V}^{\ast}(v^{\lambda}_{n})\in\mathbb{Z}^{n-1}$ from the vertex $i_{V}^{\ast}(\lambda)$ of $i_{V}^{\ast}(\Delta)$ form a $\mathbb{Z}$-basis of $\mathbb{Z}^{n-1}$.
Note that the first condition in Definition 1.2 are equivalent to the condition that the polytope $i_{V}^{\ast}(\Delta)$ is simple. In Section 5 we demonstrate examples of subspaces that are admissible and not admissible with respect to some Delzant polytopes.
@@
Let $\Delta$ be a Delzant polytope in $(\mathfrak{t}^{n})^{\ast}$, $V=\mathbb{R}p_{1}+\cdots+\mathbb{R}p_{n-1}+a$ an (n1)(n-1)-dimensional affine subspace in $\mathfrak{t}^{n}$ with rational slope, $X$ a toric manifold associated with the Delzant polytope $\Delta$. If $V$ is admissible with respect to $\Delta$, then the rank of the Jacobian matrix $Df^{\lambda}$ is equal to one at any point of the zero locus of $f^{\lambda}$ for any vertex $\lambda$ of the Delzant polytope $\Delta$. In particular, $\overline{C(V)}$ is a smooth complex hypersurface in the toric manifold $X$.
@@
If $V$ is admissible with respect to the Delzant polytope $\Delta$ of the toric manifold $X$, then the convex polytope $i_{V}^{\ast}(\Delta)$ is a Delzant polytope of the complex hypersurface $\overline{C(V)}$ in $X$.
@@
Let $V=\mathbb{R}p_{1}+\cdots+\mathbb{R}p_{n-1}+a$ be an (n1)(n-1)-dimensional affine subspace in $\mathfrak{t}^{n}$ with rational slope. If the rank of the Jacobian matrix $Df^{\lambda}$ is equal to one at any point of the zero locus of $f^{\lambda}$ for any vertex $\lambda$ of the Delzant polytope $\Delta$ of the toric manifold $X$, then $V$ is admissible with respect to $\Delta$.
@@ 1.3. Outline
This paper is organized as follows. In Section 2, we prepare for something about torus-equivariantly embedded toric hypersurfaces. In Section 3, we show Theorem 1.3. In Section 4, we show Theorem 1.5. In Section 5, we demonstrate some examples to illustrate admissible subspaces.
@@
The author is grateful to the advisor, Manabu Akaho, for a lot of encouragements and supports. The author would also like to thank Yuichi Kuno and Yasuhito Nakajima for helpful discussions, and would also like to thank the anonymous referee for helpful comments which improve the exposition. This work was supported by JST SPRING, Grant Number JPMJSP2156.
2. Review of Torus-equivariantly Embedded Toric Manifolds
In this section, we review the settings in [Yam25]. We also show some properties of the rank of the Jacobian matrix $Df^{\lambda}$ used in this paper.
@@ 2.1. Symplectic Toric Manifolds
${{}^{t}[u^{\lambda}_{1}\;\cdots\;u^{\lambda}_{n}]}[v^{\lambda}_{1}\;\cdots\;v^{\lambda}_{n}]=E_{n}$
holds for any vertex $\lambda$ of $\Delta$ (see [Yam25, Lemma 3.10] for example).
@@
The detailed construction in terms of the notation used here is given in [Yam25, Section 2.1 and Section 3.2]. Moreover, the complement of toric divisors in a $2n$-dimensional symplectic toric manifold can be identified with a complex $n$-dimensional torus.
@@ 2.2. Torus-equivariantly Embedded Toric Hypersurfaces
Let $p_{1},\ldots,p_{n-1}\in\mathbb{Z}^{n}$ be primitive vectors which are linearly independent, and $a\in\mathbb{R}^{n}$. We consider the (n1)(n-1)-dimensional affine subspace $V=\mathbb{R}p_{1}+\cdots+\mathbb{R}p_{n-1}+a$ in $\mathfrak{t}^{n}\cong\mathbb{R}^{n}$. Write $\vec{V}$ as the linear part of $V$. Assume that $V$ has a rational slope, i.e. we may assume that $p_{1},\ldots,p_{n-1}\in\mathbb{Z}^{n}$ form a $\mathbb{Z}$-basis of $\vec{V}\cap\mathbb{Z}^{n}\cong\mathbb{Z}^{n-1}$. Let $q\in\mathbb{Z}^{n}$ be a primitive basis of the orthogonal subspace to $\vec{V}$ in $\mathfrak{t}^{n}\cong\mathbb{R}^{n}$.
Through the log-affine coordinate $(\mathbb{C}^{\ast})^{n}\cong\mathfrak{t}^{n}\times T^{n}$, we define the complex subtorus $C(V)\cong V\times T^{n-1}$ (see [Yam25, Proposition 4.2] for detail). Recall that the complement of toric divisors in a toric manifold $X$ can be identified with a complex $n$-dimensional torus $(\mathbb{C}^{\ast})^{n}$. The closure $\overline{C(V)}$ of $C(V)$ in the toric manifold $X$ is expressed as follows:
@@
The closure $\overline{C(V)}$ of $C(V)$ in the toric manifold $X$ is defined as $\overline{C(V)}=\bigcup_{\lambda\in\Lambda}\varphi_{\lambda}^{-1}(\overline{C_{\lambda}(V)})$, where $\overline{C_{\lambda}(V)}\subset\varphi_{\lambda}(U_{\lambda})\cong\mathbb{C}^{n}$ is the zero locus of $f^{\lambda}$. Here, $f^{\lambda}$ are defined by
@@
Also, $\overline{C(V)}$ can be reprenseted as the zero locus of the polynomial $f$ of $X$ with $f|_{\varphi_{\lambda}(U_{\lambda})}=f^{\lambda}$. We define a singular point of $\overline{C(V)}$ in $X$ by the point where the Jacobian matrix $Df^{\lambda}$ of $f^{\lambda}$ becomes zero for some $\lambda$. In this case, we say that $\overline{C(V)}$ has a singularity. Similarly, we say that $\overline{C(V)}$ is nonsingular if $\overline{C(V)}$ has no singular point.
@@
The rank of the Jacobian matrix $Df^{\lambda}$ of $f^{\lambda}$ might be equal to zero at some points in the toric divisors. In particular, $\overline{C(V)}$ might have a singularity at some points in the toric divisors.
Note that if $\overline{C(V)}$ is a smooth complex submanifold in the toric manifold $X$, then $\overline{C(V)}$ has the effective Hamiltonian $T^{n-1}$-action through the map $i_{V}:T^{n-1}\to T^{n}$ and the $T^{n}$-action on $X$ (see [Yam25]).
@@
In this case, the rank of the Jacobian matrix $Df^{\lambda}$ is independent of the choice of $a\in\mathbb{R}^{n}$ of the affine subspace $V=\mathbb{R}p_{1}+\cdots+\mathbb{R}p_{n-1}+a$ [Yam24, Proposition 2.12], i.e. we may assume that $a=0$.